OpenStudy (anonymous):

pleasee help im trying to finish a test :D what are the first three terms of the arithmetic series if a1 = -1, an = -115, and Sn = -1160?

OpenStudy (callisto):

Are you doing your test right now?

OpenStudy (anonymous):

yess im timed

OpenStudy (anonymous):

half hr left and im stuck on 6 questions

OpenStudy (anonymous):

out of 25

OpenStudy (zzr0ck3r):

are these saposed to be three different series?

OpenStudy (callisto):

We're not supposed to help you cheat in your test. Here's something you may find it useful. 1. Sn = (a1+an) n /2 <- find n here. 2. an = a1 + (n-1)d <- find d here 3. find the three term you need using the formula in (2).

OpenStudy (anonymous):

a. -1160, -1166, -1172 c. 20, 14, 8 b. -1, -7, -13 d. -115, -121, -127

OpenStudy (callisto):

Assume Sn is the sum of AS up to nth term.

OpenStudy (anonymous):

its summer school, i just need this credit. and i dont understand what your saying

OpenStudy (callisto):

Are you allowed to get help while you're doing your test?

OpenStudy (anonymous):

im at home yeahh im allowed to use whatever i want but i cant find anything online that helps

OpenStudy (callisto):

Still, I'm not sure if it is considered as cheating, especially when it is a *test* As mentioned above, way to solve this question: 1. Sn = (a1+an) n /2 <- find n here. 2. an = a1 + (n-1)d <- find d here 3. find the three term you need using the formula in (2). Sub. the values you have, find the values you need.

OpenStudy (anonymous):

so i used these Sn = (a1+an) n /2 <- find n here. an = a1 + (n-1)d <- find d here and i got -60?

OpenStudy (anonymous):

idk wth im doing hahah

OpenStudy (callisto):

not possible.

OpenStudy (anonymous):

-1160=(-1r-115) right

OpenStudy (callisto):

no.

OpenStudy (anonymous):

its cool im just gunna gess i wasted too much time on thsi one question

OpenStudy (callisto):

Sn = (a1+an) n /2 a1 = -1 an = -115 Sn = -1160